# Optimal Execution in Intraday Energy Markets under Hawkes Processes with Transient Impact

### v1

@@ Abstract
This paper investigates optimal execution strategies in intraday energy markets through a mutually exciting Hawkes process model. Calibrated to data from the German intraday electricity market, the model effectively captures key empirical features, including intra-session volatility, distinct intraday market activity patterns, and the Samuelson effect as gate closure approaches. By integrating a transient price impact model with a bivariate Hawkes process to model the market order flow, we derive an optimal trading trajectory for energy companies managing large volumes, accounting for the specific trading patterns in these markets. A back-testing analysis compares the proposed strategy against standard benchmarks such as Time-Weighted Average Price (TWAP) and Volume-Weighted Average Price (VWAP), demonstrating substantial cost reductions across various hourly trading products in intraday energy markets.
@@ 1 Introduction
Intraday markets play a crucial role in developed energy markets, with trading volumes reaching new peaks in recent years, as observed in the German market in 2024 [21]. One of the primary drivers behind this growth is the increasing share of renewable energy, which requires market participants to balance their positions in short-term intraday markets due to the variable and uncertain production from sources like wind and solar. This behavior results in higher trading volumes and elevated price volatility as delivery time approaches, a well-documented phenomenon in commodity and energy markets known as the Samuelson effect [33].
Empirical studies, including [24] and [28], have validated the Samuelson effect in short-term energy markets. These studies highlight an exponential rise in order book activity as gate closure nears, with a surge in market and limit order arrivals driven by participants’ urgency to balance positions. Such dynamics pose significant operational challenges for energy companies, particularly when managing substantial trading volumes arising from unexpected outages, fluctuations in renewable generation, or strategic trading decisions. The central question for market participants is how to execute these trades efficiently while minimizing costs and mitigating market impact.
This paper addresses this challenge by developing optimal execution strategies for renewable energy producers in intraday energy markets. The proposed framework aims to minimize trading costs by leveraging a price impact model embedded within a limit order book setting using a mutually exciting Hawkes process to model market order flows. This study provides practical strategies for energy companies active in intraday energy markets by explicitly accounting for the unique trading activity patterns in these markets. In particular, it captures the exponential increase in trading volume approaching gate closure, followed by a sharp decline after market decoupling, where ne hour before delivery, international orders are no longer coupled with the German market, and 30 minutes before delivery, trading transitions to local, within-grid area transactions.
@@ 1.1 Stylized Facts of Price Dynamics in Intraday Energy Markets
Modeling the dynamics of intraday energy markets requires a nuanced understanding of the limit order book (LOB), particularly at high frequencies where granular trading behaviors dominate. The LOB aggregates market orders, limit orders, and cancellations, which collectively shape price movements. Intraday energy markets, being inherently high-frequency, necessitate models that explicitly capture these LOB dynamics to accurately reflect trading conditions.
@@
A distinctive feature of intraday markets is the variation in LOB activity throughout the trading session. Order flow and price changes intensify as delivery approaches, driven by increased rebalancing urgency and improved forecast accuracy. Figure 1 illustrates mid-price fluctuations for hourly delivery products at 18:00, 19:00, and 20:00 on September 30, 2023, showing higher volatility near the beginning and end of the trading sessions.
@@
Figure 1: German transaction prices for trading sessions on September 30, 2023, for hourly delivery products at 18:00, 19:00, and 20:00, relative to time to gate closure.
@@
Figure 2 highlights the empirical intensity of buy (see Figure 28 in the Appendix C for the sell side of the order book) market orders throughout the trading day, which peaks during the final hours before delivery. The midday activity increase, which is highlighted in the empirical intensity of buy (see Figure 28 for the sell side of the order book), corresponds to higher liquidity during noon delivery hours (products with a delivery 12:00-15:00), likely driven by Germany’s rapid expansion of solar energy capacity [20] as well as the general expected increase of trading activity mid-day. Conversely, activity diminishes sharply during the final trading hour due to market decoupling, which restricts cross-border trading and limits transactions to within Germany. These patterns underscore the importance of modeling the time-varying nature of LOB activity.
Figure 2: Empirical intensities of buy market orders calculated by counting the average buy events during $[t,t+\delta t]$.
Figure 3 illustrates the time-varying behavior of a lag-dependent volatility estimator along the trading horizon for all German hourly intraday products. Examining different times to maturity, we observe that volatility intensifies as delivery time approaches. From this figure, the presence of microstructure noise is evident from the increasing variance at higher frequencies (smaller $\delta$); which is evident for all the 24 trading products. This pattern suggests mean reversion at these scales. In contrast, if the price data followed a Brownian motion, the signature plot would remain flat, indicating no microstructure effects and no dependence of volatility on the time lag.
@@
This observation motivates the use of Hawkes processes for modeling price dynamics, as they capture the self-exciting and clustering behaviors characteristic of high-frequency financial data, which are properties not well-modeled by Brownian motion or simpler point processes like Poisson processes. By allowing the intensity to depend on its own path, Hawkes processes naturally account for the autocorrelation present in mid-price returns and order arrivals. Another reason to favor Hawkes processes over, e.g., time inhomogeneous Poisson processes, is that time changed arrival times of market orders cannot be adequately modeled using exponential distributions [17]. Furthermore, we observe that the integrated volatility of the mid-price process varies over the trading horizon, increasing as gate closure nears (see Figure 3). Again, this phenomenon reflects the heightened trading intensity and market activity driven by increasingly accurate renewable energy forecasts as delivery approaches. As traders receive more reliable information closer to delivery time, they adjust their positions accordingly, resulting in increased volatility.
Figure 3: Mean signature volatility surface for all hourly products with sampling frequency $\delta=\{0,\dots,300\}$ and omitting one hour to gate closure from the LOB data.
@@
These empirical observations underscore the need for models that effectively capture both the increasing intensity near gate closure and intra-session volatility patterns. In Section 2, we introduce and study the marked Hawkes price model used to describe the market order flow in detail. In Section 3, we demonstrate that the Hawkes process framework, with time-varying baseline intensities and event clustering through the self/cross excitation kernels, provides a robust approach to addressing these dynamics. Specifically, we calibrate the Hawkes model to German intraday energy market data using maximum likelihood estimation (MLE), ensuring an accurate fit to the observed market behavior, see Tables 1 and 2 below for results on the calibrated parameter space and Section B.1 for a comprehensive study on the Goodness-of-fit analysis of the different models.
@@
Moving beyond traditional parametric approaches, we employ a spline-based method with fixed knots to model baseline intensity changes over time. This non-parametric approach offers the flexibility to capture complex intensity patterns, including the exponential increase in activity as gate closure approaches. Additionally, it accommodates the sharp decline in activity observed during the final trading hour, following market decoupling [8], a phenomenon often overlooked in existing studies.
@@ 1.2 Optimal Execution in Intraday Energy Markets
Beyond modeling order flow, it is essential to evaluate how these dynamics influence trading strategies. For energy companies managing substantial trading volumes, optimal timing and execution are critical to minimizing both market impact and trading costs. By leveraging Hawkes processes to model intraday dynamics, traders can anticipate periods of high liquidity or volatility, enabling them to adjust their strategies more effectively. While much of the literature on optimal execution focuses on equity markets and price impact models [5, 22, 31, 9, 30], research on optimal execution in energy markets remains limited. Existing studies [26, 1, 23, 25] often overlook the granular effects of limit order book (LOB) events on execution prices. Furthermore, most works assume that prices evolve as continuous martingales under passive trading, incorporating all order impacts. However, at high frequencies, large orders exert substantial effects that invalidate these assumptions [2, 3].
To overcome these limitations, we employ a high frequent, tick-by-tick execution price model based on Hawkes processes, which effectively captures the influence of market orders from other participants in intraday energy markets.
High-frequency models [6, 7, 29] are well-suited for capturing key statistical properties such as autocorrelation in trade signs, volatility clustering and microstructure noise. While these approaches have primarily been applied to equity markets, recent extensions have begun to explore energy markets [17, 18], incorporating marked Hawkes processes to model mid-price jump sizes. These developments lay the foundation for more nuanced and effective optimal execution strategies tailored to the unique characteristics of energy markets.
This paper contributes to the literature on optimal execution in intraday energy markets by combining high frequency price models [17, 18] with optimal execution frameworks. Hawkes processes in our model effectively capture the flow of market orders, while price impact models are used to calculate trading costs. Additionally, our approach incorporates time-dependent baseline intensities to reflect the sharp increase in market activity near gate closure [19, 27, 24] and market decoupling thereafter. The transient market impact component, which accounts for the influence of past trades on prices, has only recently been introduced in electricity markets [25] with focus on energy arbitrage and battery energy storage system optimization problems. To the best of our knowledge, no existing work has implemented a fit of the model and an optimal trading strategy directly to real-market data, making this study a novel contribution to the field. Our backtesting analysis in Section 4 addresses the following key questions:
@@ 1.3 Structure of the Paper
This paper is structured as follows. Section 2 introduces the Hawkes price model and formulates the optimal execution problem. It presents the derivation of the optimal execution strategy using Hawkes processes to model buy and sell market orders, as detailed in Theorem 2.2, enabling a quantitative analysis of execution costs. Section 3 details the calibration protocol for the model using EPEX Spot LOB data. The Hawkes parameters and its baseline intensity are estimated independently for hourly German intraday market products, and the goodness-of-fit is evaluated with comparisons between models under varying assumptions about event dependencies and intensities. Section 4 presents the back-testing analysis, assessing the optimal execution strategy against TWAP and VWAP benchmarks using real market data from 2023–2024. The appendices include additional results on Maximum Likelihood Estimation (Section A), goodness-of-fit analysis (Section B), and various complementary findings (Section C).
@@ 2 The Hawkes Price Model and Optimal Execution Strategies
In this section, we provide an introduction to the Hawkes price model used to capture tick-by-tick price data in intraday energy markets and recall a few important properties of the model from [17]. Thereafter, we formulate the optimal execution problem for a generic model.
@@ 2.1 The Marked Hawkes Price Model
We introduce the Marked Hawkes Price (MHP) model using Poisson random measures and counting processes. Consider $M$ maturities $0<T_{1}<T_{2}<\dots<T_{M}=T$, where $M\in\mathbb{N}$. Let $K\subseteq\mathbb{R}_{+}$ denote the set of possible marks (e.g., price jumps or trade sizes), and let $\mathcal{K}$ be the Borel sigma-algebra on $K$. On a probability space $(\Omega,\mathcal{F},\mathbb{P})$, we consider $2M$ independent marked point processes representing upward and downward price jumps for each maturity.
@@
For each maturity $T_{m}$, where $m=1,\dots,M$, we define the counting processes $N_{m}^{+}(t)$ and $N_{m}^{-}(t)$, which count the number of upward and downward price jumps up to time $t$, respectively. Each jump is associated with a mark $y\in K$. These counting processes can be represented via Poisson random measures $\pi_{m}^{+}(dt,dy)$ and $\pi_{m}^{-}(dt,dy)$ on the measurable space $\left(E,\mathcal{E}\right)=\left(\mathbb{R}_{+}\times K,\mathcal{B}\left(\mathbb{R}_{+}\right)\otimes\mathcal{K}\right)$, each with intensity measure $\lambda_{m,t}^{\pm}\,dt\otimes\nu(dy)$, where $\nu$ is a probability measure on $(K,\mathcal{K})$ satisfying $\nu(\{0\})=0$. The cumulative upward and downward price jump processes for maturity $T_{m}$ are given by:
$S_{m,t}^{\pm}=\int_{0}^{t}\int_{K}y\,\pi_{m}^{\pm}(ds,dy)=\sum_{k=1}^{N_{m}^{\pm}(t)}Y_{m,k}^{\pm},$ (1)
@@
where $Y_{m,k}^{\pm}$ are the marks associated with the $k$-th upward or downward jump, respectively. The net price process is
$S_{m,t}$ $=S_{m,0}+S_{m,t}^{+}-S_{m,t}^{-}$
$=S_{m,0}+\int_{0}^{t}\int_{K}y\left(\pi_{m}^{+}(ds,dy)-\pi_{m}^{-}(ds,dy)\right),$ (2)
where $S_{m,0}$ is the initial price for maturity $T_{m}$. To incorporate clustering effects characteristic of Hawkes processes, we specify that the conditional intensities $\lambda_{m,t}^{\pm}$ of the counting processes $N_{m}^{\pm}(t)$ depend on the history of past events. Specifically, for $t\geq 0$, the conditional intensities are defined as:
$\lambda_{m,t}^{\pm}=\mu_{m}^{\pm}(t)+\sum_{n=1}^{M}\int_{0}^{t^{-}}\int_{K}\phi_{mn}^{\pm}(t-s,y)\,\pi_{n}^{\pm}(ds,dy),$ (3)
where $\mu_{m}^{\pm}(t)$ are deterministic baseline intensities, and $\phi_{mn}^{\pm}(t-s,y)$ are non-negative kernel functions representing the self-exciting and mutually exciting effects between different maturities and directions (upward or downward).
Assuming the marks are independent of the past and identically distributed according to $\nu$, we can integrate over the marks to obtain:
[…]
where the integrated kernels are defined as:
$\phi_{mn}^{\pm}(t-s)=\int_{K}\phi_{mn}^{\pm}(t-s,y)\,\nu(dy).$ (5)
In this framework, the counting processes $\{N_{m}^{\pm}(t)\}_{m=1}^{M}$ form a multivariate Hawkes process. The intensities $\lambda_{m,t}^{\pm}$ capture the dependence on past events through the convolution of the kernel functions with the counting processes, reflecting the self-exciting and mutually exciting nature of the price jumps.
In the multivariate Hawkes process, each component $i$ has a baseline intensity $\lambda_{\infty}^{i}(t)$, representing its base rate, which may vary over time. An event in component $j$ causes a jump in the intensity $\lambda^{i}(t)$ of component $i$, with the jump size determined by the decay function $\phi^{ij}(\cdot)$. The intensity $\lambda^{i}(t)$ then decays back toward its baseline according to $\phi^{ij}(\cdot)$. Self-excitation occurs when events in component $i$ affect its own intensity, while cross-excitation occurs when events in component $j$, with $i\neq j$, affect $\lambda^{i}(t)$. We model both effects.
As an example, consider a univariate Hawkes process with an exponential decay kernel, given by:
[…]
where $\lambda_{0}>0$ is the baseline intensity, $\alpha>0$ is the excitation parameter, and $\beta>0$ is the decay rate, see Figure 4.
(a) Intensity process $\lambda_{t}$
(b) Counting process $N_{t}$

### v2

@@ Abstract
This paper investigates optimal execution strategies in intraday energy markets through a mutually exciting Hawkes process model. Calibrated to data from the German intraday electricity market, the model effectively captures key empirical features, including intra-session volatility, distinct intraday market activity patterns, and the Samuelson effect as gate closure approaches. By integrating a transient price impact model with a bivariate Hawkes process to model the market order flow, we derive an optimal trading trajectory for energy companies managing large volumes, accounting for the specific trading patterns of these markets. A back-testing analysis compares the proposed strategy against standard benchmarks such as Time-Weighted Average Price (TWAP) and Volume-Weighted Average Price (VWAP), demonstrating substantial cost reductions across various hourly trading products in intraday energy markets.
@@ 1 Introduction
Intraday markets play a crucial role in developed energy markets, with trading volumes reaching new peaks in recent years, as observed in the German market in 2024 [EPEXSpotSE]. One of the primary drivers behind this growth is the increasing share of renewable energy, which requires market participants to balance their positions in short-term intraday markets due to the variable and uncertain production from sources like wind and solar. This behavior results in higher trading volumes and elevated price volatility as delivery time approaches, a well-documented phenomenon in commodity and energy markets known as the Samuelson effect [c706bac4-9c99-3ef5-b2ae-197e56e4542c].
Empirical studies, including [Luckner] and [f318e8a6-3566-30d7-a5ea-473f21c64fcd], have validated the Samuelson effect in short-term energy markets. These studies highlight an exponential rise in order book activity as gate closure nears, with a surge in market and limit order arrivals driven by participants’ urgency to balance positions. Such dynamics pose significant operational challenges for energy companies, particularly when managing substantial trading volumes arising from unexpected outages, fluctuations in renewable generation, or strategic trading decisions. The central question for market participants is how to execute these trades efficiently while minimizing costs and mitigating market impact.
This paper addresses this challenge by developing optimal execution strategies for renewable energy producers in intraday energy markets. The proposed framework aims to minimize trading costs by leveraging a price impact model embedded within a limit order book setting using a mutually exciting Hawkes process to model market order flows. This study provides practical strategies for energy companies active in intraday energy markets by explicitly accounting for the unique trading activity patterns in these markets. In particular, it captures the exponential increase in trading volume approaching gate closure, followed by a sharp decline after market decoupling, where one hour before delivery, international orders are no longer coupled with the German market, and 30 minutes before delivery, trading transitions to local, within-grid area transactions.
@@ 1.1 Stylized Facts of Price Dynamics in Intraday Energy Markets
Modeling the dynamics of intraday energy markets requires a nuanced understanding of the limit order book (LOB), particularly at high frequencies where granular trading behavior dominates. The LOB aggregates market orders, limit orders, and cancellations, which collectively shape price movements. Intraday energy markets, being inherently high-frequency, necessitate models that explicitly capture these LOB dynamics to accurately reflect trading conditions.
@@
A distinctive feature of intraday markets is the variation in LOB activity throughout the trading session. Order flow and price changes intensify as delivery approaches, driven by increased urgency to rebalance and improved forecast accuracy. Figure 1 illustrates mid-price fluctuations for hourly delivery products at 18:00, 19:00, and 20:00 on September 30, 2023, showing higher volatility near the beginning and end of the trading sessions.
@@
Figure 1: German transaction prices for trading sessions on September 30, 2023, for hourly delivery products at 18:00, 19:00, and 20:$0$, relative to time to gate closure.
@@
Figure 2(a) highlights the empirical intensity of buy (see Figure 26 in Section C for the sell side) market orders throughout the trading day, which peaks during the final hours before delivery. The midday activity increase, which is highlighted in the empirical intensity of buy orders (see Figure 26 for the sell side of the order book), corresponds to higher liquidity during noon delivery hours (products with a delivery 12:00-15:00), likely driven by Germany’s rapid expansion of solar energy capacity [Ember] as well as the general expected increase of trading activity mid-day. Conversely, activity diminishes sharply during the final trading hour due to market decoupling, which restricts cross-border trading and limits transactions to within Germany. These patterns underscore the importance of modeling the time-varying nature of LOB activity.
Figure 2(b) displays signature plots of realized volatility for all $24$ German hourly intraday products. For a given terminal horizon $T$ and sampling interval $\Delta$ (in seconds), we compute
$C_{T}(\Delta)=\sum_{k=1}^{\lfloor T/\Delta\rfloor}\!\big(S_{t_{k}}-S_{t_{k-1}}\big)^{2},\qquad t_{k}=k\,\Delta,$
where $S_{t_{k}}$ denotes the mid-price at time $t_{k}$. The figure overlays four surfaces, corresponding to $T\in\{8,7,6,5\}\,\mathrm{h}$, mapping $\Delta\mapsto C_{T}(\Delta)$ across the hourly trading products. Two main patterns emerge: Firstly, the Samuelson effect, i.e. for any fixed $\Delta$, $C_{T}(\Delta)$ increases as time to delivery shortens (from $T=8$ h to $T=5$ h), reflecting higher intraday volatility closer to gate closure in line with [Deschatre2023, DG22, RePEc:arx:papers:2211.13002]. Secondly, microstructure noise, as for each product and horizon $T$, $C_{T}(\Delta)$ decreases as $\Delta$ grows over the 0300300 second range. This pattern is consistent with noise-induced inflation of high-frequency quadratic variation and mean reversion at very short scales. If prices followed a pure Brownian motion without microstructure effects, the signature plot would remain flat in $\Delta$, showing no dependence of volatility on the sampling frequency.
@@
These empirical observations highlight the need for models that capture both the rise in volatility as gate closure approaches and the intra-session clustering of activity. Hawkes processes provide a natural framework for this task: by allowing the intensity to depend on its own history, they reproduce the self-exciting and clustering behavior characteristic of high-frequency financial data, features that Brownian motion or simpler point processes such as Poisson models cannot adequately represent. In particular, Hawkes dynamics account for the strong autocorrelation observed in mid-price returns and order arrivals in intraday energy markets. Another reason to favor Hawkes processes over, e.g., time inhomogeneous Poisson processes, is that time changed arrival times of market orders cannot be adequately modeled using exponential distributions [DG22]. Furthermore, we observe that the integrated volatility of the mid-price process varies over the trading horizon, increasing as gate closure nears (see Figure 2(b)). Again, this phenomenon reflects the heightened trading intensity and market activity driven by increasingly accurate renewable energy forecasts as delivery approaches. As traders receive more reliable information closer to delivery time, they adjust their positions accordingly, resulting in increased volatility.
(a) Empirical intensity of buy market orders (1/h) over trading time.
(b) Signature-plot surfaces $C_{T}(\Delta)$ for the 24 hourly products.
Figure 2: (a) Empirical buy-side intensity over the trading day. (b) Realized-volatility signature surfaces across sampling scales and delivery hours; layered horizons $T\in\{8,7,6,5\}$ h with x-axis the sampling interval $\Delta$ (s); y-axis the delivery hour/products from 1–24 and z-axis shows $C_{T}(\Delta)$ on $[0,T]$. Layers correspond to $T\in\{8,7,6,5\}$ h (counted back from gate closure). $C_{T}(\Delta)$ decreases in $\Delta$ (indicating microstructure noise) and, at fixed $\Delta$, rises near delivery (showing the Samuelson effect).
@@
In Section 2, we provide a concise introduction to the marked Hawkes price model used to describe the market order. Moreover, in Section 3, we demonstrate that the Hawkes process framework, with time-varying baseline intensities and event clustering through the self/cross excitation kernels, provides a robust approach to addressing these dynamics. Specifically, we calibrate the Hawkes model to German intraday energy market data using maximum likelihood estimation (MLE), ensuring an accurate fit to the observed market behavior, see Tables 1 and 2 below for results on the calibrated parameter space and Section B.1 for a comprehensive study on the goodness-of-fit analysis of the different models.
@@
Moving beyond traditional parametric approaches, we employ a spline-based method with fixed knots to model baseline intensity changes over time. This non-parametric approach offers the flexibility to capture complex intensity patterns, including the exponential increase in activity as gate closure approaches. Additionally, it accommodates the sharp decline in activity observed during the final trading hour, following market decoupling [Bal22], a phenomenon often overlooked in existing studies.
@@ 1.2 Optimal Execution in Intraday Energy Markets
Beyond modeling order flow, it is essential to evaluate how these dynamics influence trading strategies. For energy companies managing substantial trading volumes, optimal timing and execution are critical to minimizing both market impact and trading costs. By leveraging Hawkes processes to model intraday dynamics, traders can anticipate periods of high liquidity or volatility, enabling them to adjust their strategies more effectively. While much of the literature on optimal execution focuses on equity markets and price impact models [almgren2000optimal, Gatheral2013, Obizhaeva2005, Bertsimas1998, NV22], research on optimal execution in energy markets remains limited. Existing studies [KZ20, AGP16, glas2020intraday, JCP24] often overlook the granular effects of limit order book (LOB) events on execution prices. Furthermore, most works assume that prices evolve as continuous martingales under passive trading, incorporating all order impacts. However, at high frequencies, large orders exert substantial effects that invalidate these assumptions [Alfonsi2016, AlfonsiExtension].
To overcome these limitations, we employ a high frequent, tick-by-tick execution price model based on Hawkes processes, which effectively captures the influence of market orders from other participants in intraday energy markets. High-frequency models [Bacry2013, BTJF16, LN19] are well-suited for capturing key statistical properties such as autocorrelation in trade signs, volatility clustering and microstructure noise. While these approaches have primarily been applied to equity markets, recent extensions have begun to explore energy markets [DG22, Deschatre2023], incorporating marked Hawkes processes to model mid-price jump sizes. These developments lay the foundation for more nuanced and effective optimal execution strategies tailored to the unique characteristics of energy markets.
This paper contributes to the literature on optimal execution in intraday energy markets by combining high frequency price models [DG22, Deschatre2023] with optimal execution frameworks. Hawkes processes in our model effectively capture the flow of market orders, while price impact models are used to calculate trading costs. Additionally, our approach incorporates time-dependent baseline intensities to reflect the sharp increase in market activity near gate closure [DESCHATRE2021105504, KIESEL201777, Luckner] and market decoupling thereafter. The transient market impact component, which accounts for the influence of past trades on prices, has only recently been introduced in electricity markets [JCP24] with focus on energy arbitrage and battery energy storage system optimization problems. To the best of our knowledge, no existing work has implemented a fit of the model and an optimal trading strategy directly to real-market data, making this study a novel contribution to the field. Our backtesting analysis in Section 4 addresses the following key questions:
@@ 1.3 Structure of the Paper
This paper is structured as follows. Section 2 introduces the Hawkes price model and formulates the optimal execution problem. It presents the derivation of the optimal execution strategy using Hawkes processes to model buy and sell market orders, as detailed in Theorem 2.2, allowing for a quantitative analysis of execution costs. Section 3 details the calibration protocol for the model using EPEX Spot LOB data. The Hawkes parameters and its baseline intensity are estimated independently for hourly German intraday market products, and the goodness-of-fit is evaluated with comparisons between models under varying assumptions about event dependencies and intensities. Section 4 presents the back-testing analysis, assessing the optimal execution strategy against TWAP and VWAP benchmarks using real market data from 2023–2024. The appendices include additional results on maximum likelihood estimation (Section A), goodness-of-fit analysis (Section B), and some complementary findings (Section C).
@@ 2 The Hawkes Price Model and Optimal Execution Strategies
In this section, we provide an introduction to the Hawkes price model used to capture tick-by-tick price data in intraday energy markets and present a few important properties of the model. Thereafter, we formulate the optimal execution problem for this model.
@@ 2.1 The Marked Hawkes Price Model
We introduce the Marked Hawkes Price (MHP) model using Poisson random measures and counting processes. Consider $M$ maturities $0<T_{1}<T_{2}<\dots<T_{M}=T$, where $M\in\mathbb{N}$. Let $K\subseteq\mathbb{R}_{+}$ denote the set of possible marks (e.g., price jumps or trade sizes), and let $\mathcal{K}$ be the Borel sigma-algebra on $K$. On a probability space $(\Omega,\mathcal{F},\mathbb{P})$, we consider $2M$ independent marked point processes representing upward and downward price jumps for each maturity or equivalently, the arrival of buy and sell market orders.
@@
For each maturity $T_{m}$, where $m=1,\dots,M$, we define the counting processes $N_{m}^{+}(t)$ and $N_{m}^{-}(t)$, which count the number of buy and sell market orders up to time $t$, respectively, and set $N_{m}(t):=\big(N_{m}^{+}(t),N_{m}^{-}(t)\big)^{\top}$. The counting processes can be represented via Poisson random measures $\pi_{m}^{+}(dt,dy)$ and $\pi_{m}^{-}(dt,dy)$ on the measurable space $\left(E,\mathcal{E}\right)=\left(\mathbb{R}_{+}\times K,\mathcal{B}\left(\mathbb{R}_{+}\right)\otimes\mathcal{K}\right)$, each with intensity measure $\lambda_{m,t}^{\pm}\,dt\otimes\nu(dy)$, where $\nu$ is a probability measure on $(K,\mathcal{K})$ satisfying $\nu(\{0\})=0$. As before we set $\pi_{m}(dt,dy)=\left(\pi_{m}^{+}(dt,dy),\pi_{m}^{-}(dt,dy)\right)$. The cumulative upward and downward price jump processes for maturity $T_{m}$ are given by:
$S_{m,t}^{\pm}=\int_{0}^{t}\int_{K}y\,\pi^{\pm}_{m}(ds,dy)=\sum_{k=1}^{N^{\pm}_{m}(t)}Y_{m,k}^{\pm},$ (1)
@@
where $Y_{m,k}^{\pm}$ are the marks associated with the $k$-th upward or downward jump, respectively. The net price process is thus given by
$S_{m,t}$ $=S_{m,0}+S_{m,t}^{+}-S_{m,t}^{-}$ (2)
$=S_{m,0}+\int_{0}^{t}\int_{K}y\left(\pi_{m}^{+}(ds,dy)-\pi_{m}^{-}(ds,dy)\right),$
where $S_{m,0}$ is the initial price for the product with maturity $T_{m}$. Note that we do not impose positivity of $S_{m,t}$ and prices may be negative, as regularly observed in intraday electricity markets. The set of marks $K\subseteq\mathbb{R}_{+}$ refers to non-negative magnitudes (e.g., absolute tick sizes), while the sign of price changes is carried by the buy/sell decomposition via $N_{m}^{+}(t)$ and $N_{m}^{-}(t)$.
To capture the clustering of order arrivals, characteristic of Hawkes processes, we define the conditional intensities of the counting processes $N_{m}(t)$ as functions of the history of past events. More precisely, for $t\geq 0$, we define the vector of conditional intensities $\lambda_{m,t}=\big(\lambda^{+}_{m,t},\lambda^{-}_{m,t}\big)^{\top}$ as
$\lambda_{m,t}=\bar{\lambda}_{m}(t)+\int_{0}^{t}\int_{K}\phi_{m}^{\pm}(t-s,y)\,\pi_{m}(ds,dy),$ (3)
where for each hourly product $m=1,\dots,M$ we denote by
  • $\bar{\lambda}_{m}(t):=\big(\lambda_{m,\infty}^{+}(t),\lambda_{m,\infty}^{-}(t)\big)^{\top}$ the deterministic baseline intensity,
  • $\phi_{m}^{\pm}(\cdot,y)$ the positive definite kernel matrix describing the self- and cross-exciting effects between buy and sell market orders.
Thus, each past event of type $(\pm)$ (buy/sell) with mark $y$ increases the intensity $\lambda_{m,t}$ at future times $t>s$ according to the kernel $\phi_{m}^{\pm}$. The conditional intensities therefore capture dependence on past events through the convolution of the kernels with the counting processes, thereby reflecting the self-exciting and, if present, mutually exciting nature of price jumps. To illustrate the mechanism in a simple setting, Figure 3 shows a simulation of a univariate Hawkes process with exponential decay kernel, where each arrival temporarily increases the intensity, which then decays back toward its baseline. Note that under standard Hawkes stability condition (spectral radius of the kernel matrix <1<1) and finite baseline intensity, the number of jumps on any finite horizon is almost surely finite [Bacry2013].
(a) Counting process $N_{t}$
(b) Intensity process $\lambda_{t}$